OpenStudy (anonymous):

If k(2x+3)(x-1)=0 and x>1, what is "k"? (-3)/2 0 2/3 1 2

OpenStudy (anonymous):

0, because neither (2x+3) nor (x-1) can be 0 and the answer is 0

OpenStudy (anonymous):

answer is zero:)

OpenStudy (ucreategames):

i was just about to say that

OpenStudy (anonymous):

if x>1?

OpenStudy (anonymous):

k(2x+3)(x-1)=0 Divide Both sides by (2x+3)(x-1)

OpenStudy (anonymous):

k = 0/ (2x+3)(x-1) = 0

OpenStudy (anonymous):